Diffusion Morning Report 02/08/05

Ted Berg tberg at snf.stanford.edu
Tue Feb 8 06:34:18 PST 2005

Good Morning All,
    Looking pretty good this morning.  SiGe tube was changed due to 
cracked sheaths, SiH4 MFC was changed. It will need a coat and test.
Epi is down again for safety B- looking into it. Ebeam is still in High 
Vac and Paul J is working on it. Injector was changed on Tylan 1..  
Everything else looks OK for now. 

  	Date: 	2/8/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Qual testing 	 
Tylan 8-SiGe 	down 	Tube pulled and MFC changed
Tylan 9-Poly 	Up 	  	18.82/30
Tylan 10-Nitride 	Up 	  	12.56/30
Tylan 11-BPSG-LTO 	Up 	  	21.45/30
Tylan 12-FGA 	Up 	  	 
Tystar 	yellow 	  	3.73/30
Doped Poly 	GONE 	  	 
TEOS 	up 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing doped 	 
Thermco Nitride 	Down 	Waiting 	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 460 PSI

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050208/1bda52f7/attachment.html>

More information about the diffusion mailing list